If one-half a number is subtracted from five-sixths of the number, the difference is 6. Find the number

Thank You!!!

Answer by stanbon(48558)

If one-half a number is subtracted from five-sixths of the number, the difference is 6. Find the number
Let the number be "x".
EQUATION:
(5/6)x - (1/2)x = 6
(5/6)x - (3/6)x = 6
(2/6)x = 6
2x=36
x=18
Cheers,
Stan H.